le* (x+1), x<-2 wH 3. If g(x) = { , which of the following statements is/are true? (cos(max), x>-2 I. g(-2) is undefined. II.Please explain to me in detail how you got your answers, showing all steps. I need all the math help I can get.
Answer #1


(3) from the given defination g(x) is not defiend at x=-2 (2) lim g(x) = lim (e+ (x+1)) =e?(-2+1) = (3) lim g(x) = lim cos(ix